# Ready Jurist One: Benchmarking Language Agents for Legal Intelligence in Dynamic Environments

### v1

@@ 1 Introduction
Law-Eval (DASSFA 2024) Yue et al. (2024) NLP Task Rule
@@
Studies on LLM-based agents in legal contexts should focus on their ability to operate in dynamic environments, mirroring the real-world interplay of legal practitioners with evolving information and procedures. Ideal environment, e.g., court trials, relies on human participants with defined roles to ensure reliability, yet such setups are prohibitively costly and difficult to scale. Thus, developing such a system faces two key challenges: 1) Comprehensive scenarios. Real-world legal scenarios encompass diverse participants and procedural stages, e.g., advising the public, drafting documents, or issuing judgments. Sufficiently rich and representative scenarios are essential for uncovering agents’ legal sense across multiple dimensions. 2) Real-world simulation. Each scenario involves diverse roles with distinct behaviors, backgrounds, and legal perspectives, all requiring a realistic and consistent portrayal. Beyond individuals, interaction must be dynamic yet procedurally grounded.
To this end, we present J1-E NVS , an open-ended legal environment where agents engage in diverse legal scenarios, completing tasks through interaction with varied participants under procedural norms. Guided by legal experts, it curates six representative environments, categorized into three levels according to environmental complexity (role diversity, interaction demands, and procedural difficulty) as shown in Figure 1. (1) Level-I comprises knowledge questioning and legal consultation: the legal agent answers the public’s progressive questions on legal knowledge or specific cases. (2) Level-II involves complaint and defence drafting: the legal agent provides step-by-step guidance to litigants in compliance with procedural requirements, and systematically collects the necessary information to complete the drafting process. (3) Level-III consists of civil and criminal courts: within formal judicial processes, the legal agent facilitates interactions among parties, ensures procedural compliance, and produces legally valid judgments. Each environmental role is grounded in real-world legal contexts and exhibits a distinct behavioral style, ensuring diversity and authenticity.
Furthermore, we introduce J1-E VAL to evaluate agent capabilities within our constructed environments. We design task-specific and fine-grained metrics to assess task completion and procedural compliance, providing ground truth for each task and adopting rule-based or LLM-based methods. These fine-grained metrics give a deeper understanding of the skill sets required for legal task execution, and offer a realistic and reliable overview of agent capabilities, facilitating further exploration into the similarities and differences between LLM-based agents and human learning mechanisms.
@@
Contributions: (A) We introduce the first interactive and dynamic legal environment for agents. Our J1-E NVS covers six of the most representative real-world scenarios in Chinese legal practice. (B) We introduce J1-E VAL to perform fine-grained evaluation of agents across different levels of legal proficiency, ranging from trainee to lawyer to judge. (C) Experimental results reveal gaps between diverse agents and real-world legal demands, offering valuable insights to guide future progress. (D) This work introduces a new paradigm for legal intelligence, shifting from static to dynamic. Beyond evaluation, it can further be extended for data generation and reinforcement learning training.
@@ 2.1 Role Agent Setting
Setup with Real-world Legal Sources. To simulate realistic scenarios, we assign different legal elements of the same event to various roles, drawing from two sources: (1) Chinese judgment documents. For scenarios in Level II & III, Chinese court judgments encompass all the essential components of a case’s entire procedural process, e.g., parties’ details, legal disputes, facts, rulings, and laws, which align perfectly with our requirements. Specifically, we collect civil and criminal cases from judgment documents and then employ LLM to extract key elements, such as plaintiff information, claims, case details, and relevant legal provisions. (2) Chinese legal articles. For scenarios in Level I focused on progressive questioning logic and authenticity, we collect expert-drafted legal articles from the HUALV website and the Civil Trial Practice Q&A literature. As these articles are structured through stepwise questions or thematic headings, we prompt LLM to leverage this inherent structure to generate coherent topic lists and corresponding answers. Topics are classified as binary (yes/no) and open-ended questions. Details are provided in Appendix B.1.
@@
Role Agents Configuration. As shown in Figure 2, different elements are assigned to three levels of roles. (1) Type I: General Public represents persons in various occupations, such as workers and farmers, and typically lack legal knowledge and thus seek knowledge-based (Env.1) or case-specific (Env.2) consultations. They are assigned topic lists, background, and behavioral styles. (2) Type II: Specific Characters have more explicit litigation needs. They are assigned individual profiles, including personal attributes (e.g., names, genders, and addresses) and case-related details (e.g., claims, statements of defence, and evidence). Besides, they also configure behavioral styles. (3) Type III: Multi-specific Characters hold specific responsibilities within courtroom proceedings and are assigned real-world legal information corresponding to their roles. For example, the Lawyers in the civil court possess detailed case information extracted from civil judgment documents, e.g., their client information, case details, claims & defence. In criminal court scenarios sourced from criminal judgment documents, the Procurator agent is assigned to charges and opinions, while the Defendant agent is assigned case details and a stated stance toward the charges. See Appendix B.3 for details.
@@ 2.2 Multi-level Environment Construction
This level involves two participants, where the general public initiates questions, and the legal agent responds. It aims to simulate the response capabilities of the legal trainee in dynamic environments. (1) K Q . Through progressive dialogue, the legal agent resolves user confusion, moving from broad themes to specific issues. The public, assigned a topic list and occupation, continues the interaction until all topics are addressed. (2) L C . This presents a complex counseling setting grounded in a specific incident. Since legal agents lack information about the public, they actively ask for relevant details to provide better responses. Similar to KQ, the interaction continues until all question lists are resolved.
@@
This level involves two participants: an individual with a specific legal need, and a legal agent who guides the interaction to progressively collect the required information and ultimately address the litigant’s request. It aims to simulate the process by which a practicing lawyer independently initiates and completes tasks based on a given agenda. (1) C D . The legal agent gathers information from the plaintiff and formulates a complete complaint as the final objective. Guided by a predefined format, the agent autonomously plans the required content through interaction. The scenario ends once the complaint is successfully generated. (2) D D . In response to the defendant’s need for a statement of defence, the legal agent engages in in-depth dialogue with the defendant, drawing on the complaint to organize arguments and ultimately generate the defence document. The end of conditions is similar to complaint drafting.
@@ 3.1 Data Details
Following the details in Sec.2.1, we collected legal articles and documents from the year 2025 to ensure the timeliness. After systematic processing, J1-E VAL sets out 508 distinct environments, including 160 at Level I, 186 at Level II, and 192 at Level III. The detailed statistics are presented in Table 4 in Appendix. Moreover, J1-E VAL encompasses a diverse and nuanced set of legal attributes aligned with real-world legal practice. As illustrated in Figure 3, K Q and L C involve different entities and procedural topics.C D , D D and C I address various types of civil rights disputes. C R covers different categories of criminal offenses.
@@ 3.2 Evaluation Methods
  • Binary accuracy (B IN ) [0, 0.5, 1] measures accuracy in binary questions by matching method: 1 for correct, 0 for incorrect, and 0.5 for conditionally qualified answers.
@@
This stage involves the task of document generation, where the quality reflects the completeness of litigant information and evidence, the understanding of claims, and the appropriate use of document templates. Two metrics are used:
@@
This stage evaluates the quality of the final judgment and procedural compliance. Given the differences between criminal and civil judgments, we adopt distinct methods. The evaluation metrics are as follows:
  • Procedural-following score (P FS ) [0-1] evaluates completeness of procedural stage. The court contains several stages, each comprising a set of mandatory actions defined based on the legal professional’s setting. A stage is deemed complete only when all presetting actions within that stage are correctly matched.
  • Judgment score (J UD ) [0-1 or 1-10] evaluates the judgment quality. For civil court, LLM evaluates the judgment against the ground truth regarding legal points, logicality, and factual accuracy. For criminal court, we use matching to calculate the crime (C RI ) accuracy. Additionally, following Chen et al. (2019), log distance is applied to calculate the deviation span for verdict (V ER ): fines and sentences. Lower V ER indicates higher accuracy.
@@ C RI V ER
Table 2: Performances among LLM-driven legal agents on J1-E NVS , where the bolded score indicates the best performance, while the underlined score represents the second-best. Note that “-” denotes failure to accomplish the task, and lower V ER indicates higher performance.
@@ 4.2 Performance among Various Legal Agents
In this section, we use GPT-4o Achiam et al. (2023) to drive our environments, where LLM-base agents enter these to accomplish set legal tasks.
@@
To assess overall performance, we compute the average scores across all environments except V ER , and rank the models by size. As shown in Figure 5, GPT-4o achieved the best performance, showing strong legal intelligence. Qwen3-Instruct-32B demonstrates performance beyond expectations, surpassing Deepseek series. Notably, although the legal-specific LLMs perform comparably to the GPT series on existing legal benchmarks Fei et al. (2024); Yue et al. (2024), they exhibit significantly weaker performance in our setting, falling behind even smaller models. This highlights a key limitation: despite possessing legal knowledge, the absence of interactive capability hampers their effectiveness in dynamic, real-world environments. Overall, legal intelligence tends to improve with increasing model size. However, even the SOTA model fails to surpass an average score of 60, underscoring the persistent challenges that LLM agents face in handling diverse and complex legal scenarios.
@@
Table 2 reports our fine-grained metrics across levels. At Level I, both general and legal-specific models perform relatively well on K Q , indicating a solid understanding of legal knowledge. However, their performance drops on the more interactive L C task, which requires proactive engagement. Notably, even legal-specific agents built on older base can match the performance of newer LLMs. Results indicate that current agents possess legal knowledge and can apply it effectively in basic legal consultations. At Level II, LLaMA-3.3 and Deepseek-v3 achieve the highest scores on F OR for C D and D D tasks, demonstrating strong template adherence in the long legal context. For document generation quality, GPT-4o performs the best. By contrast, legal-specific and smaller models, e.g., Ministral-Instruct and LawLLM, fail to complete the assigned tasks. Detailed results of the document (D OC ) components are shown in Table 6 and 7 of Appendix.
@@
Civil and criminal courts involve distinct procedures and judgment forms. The civil court comprises five stages, yet most models struggle to complete the full process (P FS ), particularly in legal-specific models, reasoning models, and smaller models. GPT-4o and Claude-3.7 perform well in P FS , R EA and L AW . The criminal court includes three stages and more participants, where closed-source models achieve similar leading performance. Additionally, Deepseek-r1’s poor performance in executing the process highlights the limitations of reasoning models in complex environments. In summary, effective reasoning and adherence to procedural protocols within long legal contexts are essential for achieving accurate legal judgments. Detailed results of judgment components are shown in Tables 8 and 9 of Appendix.
@@
To further analyze the performance of legal agents in the civil and criminal court, we measure the completion rate by computing how many samples successfully reach designated stages. As shown in Figure 5, legal agents struggle to complete necessary actions in the Preparation stage in both civil and criminal courts. In addition, despite fewer stages and actions, navigating through the criminal court poses greater challenges for legal agents, as the environment contains more characters that increase the complexity of interaction. These findings underscore the need for enhanced procedural learning for legal agents, as well as the importance of strengthening their communication and coordination with different characters.
@@ 4.3 Can LLMs drive J1-ENVS?
Existing works Fan et al. (2025); Mou et al. (2025) have shown the strong role-playing and community modeling abilities of LLMs. In this section, we use GPT-4o and Qwen3-Instruct-32B to drive the J1-E NVS separately, demonstrating the reliability.
@@
We analyze the behavioral consistency of environment roles during interactions with three legal agents. Both GPT-4o and human evaluators rate the consistency between each role’s profile and its behavior on a scale of 1 to 10, GPT-4o evaluated all samples, while human evaluators assessed 10% of the samples per scenario. The average scores are shown in the Table 3. Across interactions with various legal agents (GPT-4o, Deepseek-v3, and Qwen3-Instruct-14B), the ratings remain consistently high and stable. Furthermore, the scores for interactions with the same agent are comparable regardless of environments. This experiment validates the reliability and effectiveness of our environment, laying a solid foundation for assessing agents’ performance in interactive legal scenarios.
@@
To further validate the stability, we employ a smaller, publicly accessible LLM (Qwen3-Instuct-32B) to drive our environment, and evaluate the overall performance of five legal agents, i.e, GPT-4o, Deepseek-V3, Qwen3-Instuct-14B, Gemma3-Instuct-14B and Chatlaw2. Figure 7 illustrates that: 1) all five agents successfully completed their respective tasks within the Qwen-based environment and achieved the same rankings as in the GPT-4o-based environment; 2) under the Qwen3-Instruct-32B environment, the performance of each agent is lower than that in the GPT-based setup, which is an acceptable outcome given the gap in model size and capability. Detailed results are shown in Table 5. Overall, the relative performance differences and rankings among agents remain consistent across various environments, further validating the robustness of our framework and highlighting its potential for deployment across diverse scenarios.
@@
We analyze the reliability of the LLM-based evaluation of various metrics, such as N BIN , D OC , J UD , and R EA . We select 10 cases per scenario. Human evaluators are asked to evaluate the performance of legal agents, and the difference between the human score and the corresponding GPT-4o score is plotted in Figure 7. We find that the scores provided by GPT-4o and human evaluators largely(over 76%) remain consistent. These results demonstrate the reliability and robustness of our evaluation framework, which enhances the validity of the findings of the legal agents.

### v2

@@ 1 Introduction
LexEval (NeurIPS 2024) Li et al. (2024) NLP Task Rule
@@
Studies on LLM-based agents in legal contexts should focus on their ability to operate in dynamic environments, mirroring the real-world interplay of legal practitioners with evolving information and procedures. The ideal environment, e.g., court trials, relies on human participants with defined roles to ensure reliability, yet such setups are prohibitively costly and difficult to scale. Thus, developing such a system faces two key challenges: 1) Comprehensive scenarios. Real-world legal scenarios encompass diverse participants and procedural stages, e.g., advising the public, drafting documents, or issuing judgments. Sufficiently rich and representative scenarios are essential for uncovering agents’ legal sense across multiple dimensions. 2) Real-world simulation. Each scenario involves diverse roles with distinct behaviors, backgrounds, and legal perspectives, all requiring a realistic and consistent portrayal. Beyond individuals, the interaction must be dynamic and yet procedurally grounded.
To this end, we present J1-E NVS , an open-ended legal environment where agents engage in diverse legal scenarios, completing tasks through interaction with various participants under procedural norms. Guided by legal experts, it curates six representative environments, categorized into three levels according to environmental complexity (role diversity, interaction demands, and procedural difficulty) as shown in Figure 1. (1) Level-I comprises knowledge questioning and legal consultation: the legal agent answers the public’s progressive questions on legal knowledge or specific cases. (2) Level-II involves complaint and defence drafting: the legal agent provides step-by-step guidance to litigants in compliance with procedural requirements, and systematically collects the necessary information to complete the drafting process. (3) Level-III consists of civil and criminal courts: within formal judicial processes, the legal agent facilitates interactions among parties, ensures procedural compliance, and produces legally valid judgments. Each environmental role is grounded in real-world legal contexts and exhibits a distinct behavioral style, ensuring diversity and authenticity.
Furthermore, we introduce J1-E VAL to evaluate the agent’s capabilities within our constructed environments. We design task-specific and fine-grained metrics to assess task completion and procedural compliance, providing ground truth for each task and adopting rule-based or LLM-based methods. These fine-grained metrics give a deeper understanding of the skill sets required for legal task execution, and offer a realistic and reliable overview of agent capabilities, facilitating further exploration into the similarities and differences between LLM-based agents and human learning mechanisms.
@@
Contributions: (A) We introduce the first interactive and dynamic legal environment for agents. Our J1-E NVS covers six of the most representative real-world scenarios in Chinese legal practice. (B) We introduce J1-E VAL to perform fine-grained evaluations of agents across different levels of legal proficiency, ranging from trainee to lawyer to judge. (C) Experimental results reveal the gaps between diverse agents and real-world legal demands, offering valuable insights to guide future progress. (D) This work introduces a new paradigm for legal intelligence, shifting from static to dynamic. Beyond evaluation, it can be extended to data generation and reinforcement learning training.
@@ 2.1 Role Agent Setting
Setup with Real-world Legal Sources. To simulate realistic scenarios, we assign different legal elements of the same event to various roles, drawing from two sources: (1) Chinese judgment documents. For scenarios in Level II & III, Chinese court judgments encompass all the essential components of a case’s entire procedural process, e.g., parties’ details, legal disputes, facts, rulings, and laws, which perfectly align with our requirements. Specifically, we collect civil and criminal cases from judgment documents and then employ LLM to extract key elements, such as plaintiff information, claims, case details, and relevant legal provisions. (2) Chinese legal articles. For scenarios in Level I focused on progressive questioning logic and authenticity, we collect expert-drafted legal articles from the HUALV website and the Civil Trial Practice Q&A literature. As these articles are structured through stepwise questions or thematic headings, we prompt an LLM to leverage this inherent structure to generate coherent topic lists and corresponding answers. Topics are classified as binary (yes/no) and open-ended questions. Details are provided in Appendix B.1.
@@
Role Agents Configuration. As shown in Figure 2, different elements are assigned to three levels of roles. (1) Type I: General Public represents persons in various occupations, such as workers and farmers, and typically lack legal knowledge and thus seek knowledge-based (Env.1) or case-specific (Env.2) consultations. They are assigned topic lists, background, and behavioral styles. (2) Type II: Specific Characters have more explicit litigation needs. They are assigned individual profiles, including personal attributes (e.g., names, genders, and addresses) and case-related details (e.g., claims, statements of defence, and evidence). Besides, they also configure behavioral styles. (3) Type III: Multi-specific Characters hold specific responsibilities within courtroom proceedings and are assigned real-world legal information corresponding to their roles. For example, Lawyers in the civil court possess detailed case information extracted from civil judgment documents, e.g., their client information, case details, claims & defence. In criminal court scenarios sourced from criminal judgment documents, the Procurator agent is assigned to charges and opinions, while the Defendant agent is assigned case details and a stated stance toward the charges. See Appendix B.3 for details.
@@ 2.2 Multi-level Environment Construction
This level involves two participants, where the general public initiates questions, and the legal agent responds. It aims to simulate the response capabilities of the legal trainee in dynamic environments. (1) K Q . Through progressive dialogue, the legal agent resolves user confusion, moving from broad themes to specific issues. The public, assigned a topic list and occupation, continues the interaction until all topics are addressed. (2) L C . This presents a complex counseling setting grounded in a specific incident. Since legal agents lack information about the public, they actively ask for relevant details to provide better responses. Similarly to KQ, the interaction continues until all questions are resolved.
@@
This level involves two participants: an individual with a specific legal need, and a legal agent who guides the interaction to progressively collect the required information and ultimately address the litigant’s request. It aims to simulate the process by which a practicing lawyer independently initiates and completes tasks based on a given agenda. (1) C D . The legal agent gathers information from the plaintiff and formulates a complete complaint as the final objective. Guided by a predefined format, the agent autonomously plans the required content through interaction. The scenario ends once the complaint is successfully generated. (2) D D . In response to the defendant’s need for a statement of defence, the legal agent engages in an in-depth dialogue with the defendant, drawing on the complaint to organize arguments and ultimately generate the defence document. The end of conditions is similar to complaint drafting.
@@ 3.1 Data Details
Following the details in Sec.2.1, we collected legal articles and documents from the year 2025 to ensure timeliness. After systematic processing, J1-E VAL sets out 508 distinct environments, including 160 at Level I, 186 at Level II, and 192 at Level III. Detailed statistics are presented in Table 4 in Appendix. Moreover, J1-E VAL encompasses a diverse and nuanced set of legal attributes aligned with real-world legal practice. As illustrated in Figure 3, K Q and L C involve different entities and procedural topics.C D , D D and C I address various types of civil rights disputes. C R covers different categories of criminal offenses.
@@ 3.2 Evaluation Methods
  • Binary accuracy (B IN ) [0, 0.5, 1] measures accuracy in binary questions using a matching method: 1 for correct answers, 0 for incorrect, and 0.5 for conditionally qualified answers.
@@
This stage involves the task of document generation, where quality reflects the completeness of litigant information and evidence, the understanding of the claims, and the appropriate use of document templates. Two metrics are used:
@@
This stage evaluates the quality of the final judgment and the procedural compliance. Given the differences between criminal and civil judgments, we adopt distinct methods. The evaluation metrics are as follows:
  • Procedural-following score (P FS ) [0-1] evaluates the completeness of the procedural stage. The court contains several stages, each comprising a set of mandatory actions defined based on the legal professional’s setting. A stage is deemed complete only when all pre-setting actions within that stage are correctly matched.
  • Judgment score (J UD ) [0-1 or 1-10] evaluates the judgment quality. For the civil court, LLM evaluates the judgment against the ground truth regarding legal points, logicality, and factual accuracy. For the criminal court, we use matching to calculate the crime (C RI ) accuracy. Additionally, following Chen et al. (2019), log distance is applied to calculate the deviation span for verdict (V ER ): fines and sentences. Lower V ER indicates higher accuracy.
@@ C RI V ER
Table 2: performance among LLM-driven legal agents on J1-E NVS , where the bolded score indicates the best performance, while the underlined score represents the second-best. Note that “-” denotes failure to accomplish the task, and lower V ER indicates higher performance.
@@ 4.2 Performance among Various Legal Agents
In this section, we use GPT-4o Achiam et al. (2023) to drive our environments, where LLM-based agents enter to accomplish set legal tasks.
@@
To assess overall performance, we compute the average scores across all environments except V ER , and rank the models by size. As shown in Figure 5, GPT-4o achieves the best performance, showing strong legal intelligence. Qwen3-Instruct-32B demonstrates performance beyond expectations, surpassing Deepseek series. Notably, although the legal-specific LLMs perform comparably to GPT series on existing legal benchmarks Fei et al. (2024); Yue et al. (2024), they exhibit significantly weaker performance in our setting, falling behind even smaller models. This highlights a key limitation: despite possessing legal knowledge, the absence of interactive capability hampers their effectiveness in dynamic, real-world environments. Overall, legal intelligence tends to improve with increasing model size. However, even the SOTA model fails to surpass an average score of 60, underscoring the persistent challenges that LLM agents face in handling diverse and complex legal scenarios.
@@
Table 2 reports our fine-grained metrics across levels. At Level I, both general and legal-specific models perform relatively well on K Q , indicating a solid understanding of legal knowledge. However, their performance drops on the more interactive L C task, which requires proactive engagement. Notably, even legal-specific agents built on an older base can match the performance of newer LLMs. Results indicate that current agents possess legal knowledge and can apply it effectively in basic legal consultations. At Level II, LLaMA-3.3 and Deepseek-v3 achieve the highest scores on F OR for C D and D D tasks, demonstrating strong template adherence in the long legal context. For document generation quality, GPT-4o performs the best. By contrast, legal-specific and smaller models, e.g., Ministral-Instruct and LawLLM, fail to complete the assigned tasks. Detailed results of the document (D OC ) components are shown in Table 6 and 7 of Appendix.
@@
Civil and criminal courts involve distinct procedures and judgment forms. The civil court comprises five stages, yet most models struggle to complete the full process (P FS ), particularly in legal-specific models, reasoning models, and smaller models. GPT-4o and Claude-3.7 perform well in P FS , R EA and L AW . The criminal court includes three stages and more participants, where closed-source models achieve similar leading performance. Additionally, Deepseek-r1’s poor performance in executing the process highlights the limitations of reasoning models in complex environments. In summary, effective reasoning and adherence to procedural protocols within long legal contexts are essential for achieving accurate legal judgments. Detailed results of the judgment components are shown in Tables 8 and 9 of Appendix.
@@
To further analyze the performance of legal agents in the civil and criminal court, we measure the completion rate by computing how many samples successfully reach designated stages. As shown in Figure 5, legal agents struggle to complete necessary actions in the Preparation stage in both civil and criminal courts. In addition, despite fewer stages and actions, navigating the criminal court poses greater challenges for legal agents, as the environment contains more characters that increase the complexity of the interaction. These findings underscore the need for enhanced procedural learning for legal agents, as well as the importance of strengthening their communication and coordination with different characters.
@@ 4.3 Can LLMs drive J1-ENVS?
Existing works Fan et al. (2025); Mou et al. (2025) have shown the strong role-playing and community modeling abilities of LLMs. In this section, we use GPT-4o and Qwen3-Instruct-32B to drive the J1-E NVS separately, demonstrating reliability.
@@
We analyze the behavioral consistency of environment roles during interactions with three legal agents. Both GPT-4o and human evaluators rate the consistency between each role’s profile and its behavior on a scale of 1 to 10, GPT-4o evaluates all samples, while human evaluators assesse 10% of the samples per scenario. The average scores are shown in Table 3. Across interactions with various legal agents (GPT-4o, Deepseek-v3, and Qwen3-Instruct-14B), the ratings remain consistently high and stable. Furthermore, the scores for interactions with the same agent are comparable regardless of the environment. This experiment validates the reliability and effectiveness of our environment, laying a solid foundation for assessing agents’ performance in interactive legal scenarios.
@@
To further validate stability, we employ a smaller, publicly accessible LLM (Qwen3-Instruct-32B) to drive our environment, and evaluate the overall performance of five legal agents, i.e, GPT-4o, Deepseek-V3, Qwen3-Instruct-14B, Gemma3-Instruct-14B and Chatlaw2. Figure 7 illustrates that: 1) all five agents successfully completed their respective tasks within the Qwen-based environment and achieves the same rankings as in the GPT-4o-based environment; 2) under the Qwen3-Instruct-32B environment, the performance of each agent is lower than that in the GPT-based setup, which is an acceptable outcome given the gap in model size and capability. Detailed results are shown in Table 5. Overall, the relative performance differences and rankings among agents remain consistent across various environments, further validating the robustness of our framework and highlighting its potential for deployment across diverse scenarios.
@@
We analyze the reliability of the LLM-based evaluation of various metrics, such as N BIN , D OC , J UD , and R EA . We select 10 cases per scenario. Human evaluators are asked to evaluate the performance of legal agents, and the difference between the human score and the corresponding GPT-4o score is plotted in Figure 7. We find that the scores provided by GPT-4o and human evaluators largely remain consistent(over 76%). These results demonstrate the reliability and robustness of our evaluation framework, which enhances the validity of the findings of the legal agents.